OpenStudy (anonymous):

Given f(x) = 2x^2 + 2x and g(x) = x + 1, find the function (f/g)(x)

OpenStudy (anonymous):

plug g(x) into the x of your f(x) function. 2(x+1)^2+2(x+1) and just solve.

OpenStudy (anonymous):

thats the part i dont get

OpenStudy (anonymous):

Do the exponent first so (x+1)^2= x^2+2x+1 then you multiply (x^2+2x+1) * 2 =2x^2+4x+2 you multiply the second part of the function 2(x+1) = 2x+2 add 2x^2+4x+2+2x+2 and your answer is : 2x^2+6x+4.